# BoxFusion: Reconstruction-Free Open-Vocabulary 3D Object Detection via Real-Time Multi-View Box Fusion

### v2

@@ 3.3.2 Correspondence Association
While the spatial association can reliably associate room-level common objects, small objects are challenging to associate since proposals for the same object are less accurate and likely to be spatially adjacent but not overlapping. Inspired by [34, 16], we argue that image-based correspondences are efficient and robust to match box proposals from different camera views that belong to the same object instance. However, directly using pre-trained models to find correspondence between the current image $I_{t}$ and all historical frames that have visibility overlap is time-consuming and computationally expensive. Therefore, we propose a simple yet effective solution that projects the global 3D boxes $\mathcal{G}_{t-1}$ at frame $t-1$ to the current image $I_{t}$, and determines whether the current candidate box $\mathcal{B}_{i}$ corresponds to any global box in $\mathcal{G}_{t-1}$ with $M$ boxes in total by the IoU of projective bounding boxes as shown in Eq. 3.
$j_{max}=arg\,max(\Phi(\mathcal{P}(T^{-1}_{t},\mathcal{B}_{i}),\mathcal{P}(T^{-1}_{t},\mathcal{G}^{j}_{t}))),j\in M,$ (3)
where $\Phi$ denotes the 2D IoU of two convex hulls obtained from the projective corners of 3D bounding boxes. $\mathcal{P}$ refers to the projection using the pin-hole camera assumption. If the IoU $\Phi_{i}{j_{max}}$ between the current candidate $\mathcal{B}_{i}$ and global box $\mathcal{G}^{j_{max}}_{t}$ is larger than $\tau_{2d}$, similar strategies in the spatial association are leveraged to determine whether $\mathcal{B}_{i}$ should be added to the candidate list $\Psi_{j_{max}}$ for further multi-view box fusion.
@@ 3.4 Multi-view Box Fusion
Since the proposals from different camera views are local and not entirely accurate, it is of great necessity to fuse these candidate boxes belonging to the same object into a multi-view consistent bounding box. For example, the sequential camera views only cover parts of a big cloth cabinet, and there would be several separate 3D boxes for the cabinet without box fusion.
As shown in Figure 3, for global 3D boxes $\mathcal{G}^{i}_{t}$ and corresponding candidate box list $\Psi^{i}_{t}$, the goal is to aggregate the associated 3D boxes in $\Psi^{i}_{t}$ and fuse them into a global one that is consistent in multiple views of $\Psi^{i}_{t}$. Empirically, we find the proposals generated by foundation models like Cubify Anything [14] look accurate in 2D images but are biased from the ground truth in 3D with scale uncertainty. Therefore, we propose to optimize the position and scales of the global box $\mathcal{G}^{i}_{t}$ to fit the proposals $\Psi^{i}_{t}$ in different views of the same object. Given the position $\mathbf{p}=(x,y,z)$ and shape $\mathbf{s}=(l,w,h)$ of $\mathcal{G}^{i}_{t}$, which are initialized as the average of those of boxes in $\Psi^{i}_{t}$, we aim to optimize the 6D variables $(\mathbf{p},\mathbf{s})=(x,y,z,l,w,h)$ to enable the global box $\mathcal{G}^{i}_{t}$ to exactly cover each candidate box in the corresponding view.
Figure 4: 2D example of particle filtering optimization (PFO) using pre-sampled swarm template (PST) for box fusion. At step $k$, particles in $\mathcal{R}^{*}_{k}$ are selected as the superior set, moving the center $q^{*}_{k-1}$ of particles $\mathcal{R}_{k}$ to $q^{*}_{k}$ and rescaling $\mathcal{R}_{k}$ to $\mathcal{R}_{k+1}$. Similar moving and rescaling is employed to the particle set $\mathcal{R}_{k+1}$ at step $k+1$. In this way, the optimization of box fusion can converge efficiently and accurately.
Let us define the objective function first. The global box $\mathcal{G}^{i}_{t}$ with parametes $(\mathbf{p},\mathbf{s})$ to be optimized is projected to images $\{I_{j}|j\in\varphi^{i}_{t}\}$ with camera poses $\{T_{j}|j\in\varphi^{i}_{t}\}$, where $\varphi^{i}_{t}$ is the set of camera indices of $\Psi^{i}_{t}$. The projected box corners of the global box $\mathcal{G}^{i}_{t}$ in image $I_{j}$ are used to derive a 2D convex hull denoted as $\alpha_{ij}$. Similarly, we can obtain the candidate box 2D convex hull $\beta_{ij}$. The goal is to maximize the 2D IoU between $\alpha_{ij}$ and $\beta_{ij}$ in each view. $\alpha_{ij}$ can be obtained by projecting the corners of $\mathcal{G}^{i}_{t}$ to the image $I_{j}$ using the camera pose $T_{j}$. The optimization can be formulated as follows:
$\alpha_{ij}=\mathcal{P}(T^{-1}_{t},(\mathcal{G}^{i}_{t}(\mathbf{p},\mathbf{s}))).$ (4)
@@
The goal is to optimize the objective function in Eq. 5, but the optimization is highly non-linear and challenging to solve in real time. Inspired by [8, 44], to address this, we propose to use a pre-sampled particle swarm template (PST) to guide the optimization. This process is based on particle filtering optimization (PFO), and the likelihood function corresponding to Eq. 5 can be formulated as 6.
@@
The PST is a set of $N_{pst}$ pre-sampled particles uniformly sampled in 6D space corresponding to the position $\mathbf{p}=(x,y,z),\mathbf{p}\in[-1,1]$ and shape $\mathbf{s}=(l,w,h),\mathbf{s}\in[0,1]$. With the initialized global box $\mathcal{G}^{i}_{t}(\mathbf{p}^{\prime},\mathbf{s}^{\prime})$, we add the PST to $\mathcal{G}^{i}_{t}(\mathbf{p}^{\prime},\mathbf{s}^{\prime})$ and detect which particle fits better using the objective function in Eq. 5. As shown in Figure 4, we select the best particles in $\mathcal{R}_{k}$ as a superior set $\mathcal{R}^{*}_{k}$, which is moved and rescaled according to the superior set $\mathcal{R}^{*}_{k}$ in step $k$. Then we can obtain the best transformation $q^{*}_{k}=(\mathbf{p}^{*}_{k},\mathbf{s}^{*}_{k})$, which are added to the current observation: $q^{\prime}=(\mathbf{p}^{\prime},\mathbf{s}^{\prime})+(\mathbf{p}^{*}_{k},\mathbf{s}^{*}_{k})$. Once the iteration reaches the maximum number or the optimization converges, we use the current observation $(\mathbf{p}^{\prime},\mathbf{s}^{\prime})$ as the optimized one. Subsequently, we update the global box $\mathcal{G}^{i}_{t}$ with the finally optimized parameters $q^{\prime}=(\mathbf{p}^{\prime},\mathbf{s}^{\prime})$. The proposed random optimization is performed once the number of candidate boxes $\Psi^{i}_{t}$ is larger than a threshold $\tau_{box}$, or new candidates are added to the list.
@@ 4.1 Experimental Setup
Baselines: We mainly compare our methods to state-of-the-art offline point-cloud based methods: FCAF [28], TR3D [29] and SpatialLM [32]. Since there are no online open-vocabulary 3D detection alternatives, we take online perception methods: EmbodiedSAM [37] and OnlineAnySeg [31] into consideration, which are initially designed for instance segmentation and are post-processed for gravity-aware 3D bounding box detection.
Metrics: We use the standard metrics for 3D object detection following [24, 1], including Average Precision (AP) at different IoU thresholds. For CA-1M and ScanNetV2, we report class-agnostic $AP$ at IoU thresholds of 0.15, 0.25, and 0.5. ScanNetV2 only provides ground-truth axis-aligned bounding boxes, which requires results of all methods are axis-aligned for fair comparison. The running efficiency of the system FPS and GPU memory usage are considered as well.
@@ 4.2 Quantitative Results
The quantitative comparison of our method and state-of-the-art baselines on CA-1M and ScanNetV2 is shown in Table 1. All sequences of the CA-1M validation set and eight sequences of ScanNetV2 are considered. Class-agnostic detection is adopted, since it is easier to obtain the categories of objects with the help of visual foundation models. Our method achieves the best performance on both datasets among the online methods, especially on CA-1M, where our method outperforms the second-best method by a large margin ($22.05$ improvement in $AP_{15}$). Point-based methods [28, 29, 32] demonstrate poor generalization on CA-1M, which is significantly lower than online alternatives in terms of $AP_{15}$ to $AP_{50}$. The main reason is that point-based methods are trained on closed-set datasets, which focus on room-level object detection and typically ignore other ordinary and daily objects.
@@
Online methods like EmbodiedSAM [37] and OnlineAnySeg [31] are designed for instance segmentation, and the results are transformed into bounding boxes for evaluation, following [15]. Details about the transformation are demonstrated in the supplementary materials. Online methods aim to detect objects with streaming RGB-D inputs, which overall shows better performance than offline methods. This may be attributed to that all online methods in Table 1 can efficiently capture the rich and useful contexts in images with VFMs for more comprehensive scene understanding.
@@
Point-based method [28, 29] outperforms the others on ScanNetV2. These methods are trained on closed-set datasets and struggle to detect objects of other common categories, which is discussed in Section 4.3. $AP_{15}$ of our method is over 30 on both CA-1M and ScanNetV2, demonstrating the robustness of our method. Moreover, the significant improvement, 20.47 and 9.55 in $AP_{25}$ compared to the best online method, illustrates that our method can detect objects effectively and accurately without the need for dense reconstruction, which is a significant advantage over other methods. Note that the FPS of OnlineAnySeg [31] is only for the fusion stage, and the segmentation of images is pre-processed, which is time-consuming. Our FPS is over 20 on average on the two datasets, and can detect and obtain the open-vocabulary semantic features or categories in real time.
Figure 5: Gallery of 3D object detection on CA-1M and ScanNetV2. Our method is reconstruction-free, and the mesh is semi-transparent. Our method can accurately detect both the common objects and the objects that are not included in most closed-set datasets. Offline methods show poor generalization on fine-grained objects, and OnlineAnySeg (represented as the online method) is slightly better than offline methods but still has a high rate of missed detection. Best view on screen.
@@
4.3 Qualitative Results
As shown in Figure 5, we compare the detections of our methods to other state-of-the-art baselines. The background of our method is semi-transparent, since we do not rely on reconstruction for detection, which is adopted for all the other methods. On the first two rows, a comparison of two sequences of the CA-1M dataset is displayed, where the ground-truth is more fine-grained than ScanNetV2. Our method stands out as the best with detailed and accurate object detections. FCAF [28] is trained on closed-set datasets, which can detect the common objects (e.g., sofas and tables) accurately but struggle to detect the fine-grained objects (e.g., newspapers on the coffee table). SpatialLM [32] can only detect the big objects accurately, and there is a significantly high rate of missed detection. OnlineAnySeg [31] can successfully detect both the common objects and previously unseen objects thanks to the VFM. However, it is hard for OnlineAnySeg to detect the small but important objects (e.g., books and keys at the upper right of the first row). This can be attributed to the merging strategies that rely on point clouds for spatial overlaps, which may fail due to the noise or error in the reconstruction. In contrast, our method is reconstruction-free and leverages the rich context in multi-view images to detect and fuse the bounding boxes of 3D objects, leading to fine-grained and accurate 3D object detection.
As shown in Figure 6, our method is open-vocabulary and supports object retrieval with user-specific text prompts. For example, if the text prompt indicates it needs to charge the phone (first row in Figure 6), the power outlet on the wall marked by the red rectangle is highlighted to correspond to the prompt. This means that our method can be easily adapted to many downstream embodied tasks.
@@ 4.3 Qualitative Results
Method $AP_{15}$uparrow $AP_{25}$uparrow FPS $\uparrow$ GPU mem. $\downarrow$
@@
4.4 Ablation Studies
  • SA: Without the SA module, the $AP_{15}$ drops by 18.64, which indicates that the spatial association is important for global 3D object detection. It can provide spatially overlapping details to associate box proposals from different views.
  • MVF: The proposed multi-view box fusion is utilized to generate consistent 3D bounding boxes based on the association modules, leading to significant improvement in $AP_{15}$ and $AP_{25}$ compared to not using the MVF module.

### v3

@@ 3.3.2 Correspondence Association
While the spatial association can reliably associate room-level common objects, it is challenging for small objects. Proposals for the same small object are typically less accurate and likely to be spatially adjacent but not overlapping. Inspired by [41, 21], we argue that image-based correspondences are efficient and robust to match box proposals from different camera views that belong to the same object instance. However, directly using pre-trained models to find correspondence between the current image $I_{t}$ and all historical frames that have visibility overlap is time-consuming and computationally expensive. Therefore, we propose a simple yet effective solution that utilizes 2D projective IoU to associate the remaining small objects. Taking the local object box $B_{i}$ in frame $t$ as an example, global 3D boxes $\mathcal{G}_{t-1}$ at frame t1t-1 are projected to the current image $I_{t}$. We check whether the local candidate box $\mathcal{B}_{i}$ corresponds to any global box in $\mathcal{G}_{t-1}$ with $K$ boxes by the IoU of projective bounding boxes as shown in Eq. 3.
$j_{max}=arg\,max(\Phi(\mathcal{P}(T^{-1}_{t},{B}_{i}),\mathcal{P}(T^{-1}_{t},\mathcal{G}^{j}_{t}))),j\in K,$ (3)
where $\Phi$ denotes the 2D IoU of two convex hulls obtained from the projective corners of 3D bounding boxes. $\mathcal{P}$ refers to the projection using the pinhole camera assumption. If the IoU $\Phi_{i{j_{max}}}$ between the current candidate $B_{i}$ and global box $\mathcal{G}^{j_{max}}_{t}$ is larger than $\tau_{2d}$, similar strategies in the spatial association are leveraged to determine whether $B_{i}$ should be added to the candidate list $\Psi^{j_{max}}_{t}$ of $\mathcal{G}^{j_{max}}_{t}$ for further multi-view box fusion.
@@ 3.4 Multi-view Box Fusion
Since the proposals from different camera views are local and not entirely accurate, it is of great necessity to fuse these candidate boxes belonging to the same object into a multi-view consistent bounding box. For example, the sequential camera views only cover parts of a big cabinet, and there would be several separate 3D boxes for the cabinet without box fusion.
As shown in Figure 3, for global 3D boxes $\mathcal{G}^{i}_{t}$ and corresponding candidate box list $\Psi^{i}_{t}$, the goal is to aggregate the associated 3D boxes in $\Psi^{i}_{t}$ and fuse them into a global one that is consistent in multiple views. Empirically, we find the box proposals generated by foundation models like Cubify Anything [19] look accurate in 2D images but are biased from the ground truth in 3D with scale uncertainty. Therefore, we propose to optimize the position and scales of the global box $\mathcal{G}^{i}_{t}$ to fit the proposals $\Psi^{i}_{t}$ in different views of the same object. Given the position $\mathbf{p}=(x,y,z)$ and shape $\mathbf{s}=(l,w,h)$ of $\mathcal{G}^{i}_{t}$, which are initialized as the average of those of boxes in $\Psi^{i}_{t}$, we aim to optimize the 6D variables $(\mathbf{p},\mathbf{s})=(x,y,z,l,w,h)$ to enable the global box $\mathcal{G}^{i}_{t}$ to exactly cover each candidate box in the corresponding views.
Figure 4: 2D example of particle filtering optimization (PFO) using pre-sampled swarm template (PST) for box fusion. At step $k$, particles in $\mathcal{R}^{*}_{k}$ are selected as the superior set, moving the center $q^{*}_{k-1}$ of particles $\mathcal{R}_{k}$ to $q^{*}_{k}$ and rescaling $\mathcal{R}_{k}$ to $\mathcal{R}_{k+1}$. Similar moving and rescaling is employed to the particle set $\mathcal{R}_{k+1}$ at step $k+1$. Therefore, the optimization can converge efficiently and accurately.
Let us define the objective function first. The global box $\mathcal{G}^{i}_{t}$ with parametes $(\mathbf{p},\mathbf{s})$ to be optimized is projected to images $\{I_{j}|j\in\varphi^{i}_{t}\}$ with camera poses $\{T_{j}|j\in\varphi^{i}_{t}\}$, where $\varphi^{i}_{t}$ is the set of camera indices of $\Psi^{i}_{t}$. The projected box corners of the global box $\mathcal{G}^{i}_{t}$ in image $I_{j}$ are used to derive a 2D convex hull denoted as $\alpha_{ij}$. Similarly, we can obtain the 2D convex hull $\beta_{ij}$ of the candidate box ($j\in\varphi^{i}_{t}$). The goal is to maximize the 2D IoU between $\alpha_{ij}$ and $\beta_{ij}$ in each view. $\alpha_{ij}$ can be obtained by projecting the corners of $\mathcal{G}^{i}_{t}$ to the image $I_{j}$ using the camera pose $T_{j}$ and $\mathcal{Q}$ for the computation of convex hulls as Eq. 4. The optimization can be formulated as follows:
$\alpha_{ij}=\mathcal{Q}(\mathcal{P}(T^{-1}_{t},(\mathcal{G}^{i}_{t}(\mathbf{p},\mathbf{s})))).$ (4)
@@
The goal is to optimize the objective function in Eq. 5, but the optimization is highly non-linear and challenging to solve in real time. Inspired by [12, 51, 37, 18], to address this, we propose to use a pre-sampled particle swarm template (PST) to guide the optimization. This process is based on particle filtering optimization (PFO), and the likelihood function corresponding to Eq. 5 can be formulated as 6.
@@
The PST is a set of $N_{pst}$ pre-sampled particles uniformly sampled in 6D space corresponding to the position $\mathbf{p}=(x,y,z),\mathbf{p}\in[-1,1]$ and shape $\mathbf{s}=(l,w,h),\mathbf{s}\in[-1,1]$. With the initialized global box $\mathcal{G}^{i}_{t}(\mathbf{p}^{\prime},\mathbf{s}^{\prime})$, we apply the PST to $\mathcal{G}^{i}_{t}(\mathbf{p}^{\prime},\mathbf{s}^{\prime})$ and detect which particle fits better using the objective function in Eq. 5. As shown in Figure 4, we select the best particles in the PST $\mathcal{R}_{k}$ as a superior set $\mathcal{R}^{*}_{k}$, which is moved and rescaled according to the superior set $\mathcal{R}^{*}_{k}$ in step $k$. Then we can obtain the best transformation $q^{*}_{k}=(\mathbf{p}^{*}_{k},\mathbf{s}^{*}_{k})$, which are added to the current observation: $q^{\prime}=(\mathbf{p}^{\prime},\mathbf{s}^{\prime})+(\mathbf{p}^{*}_{k},\mathbf{s}^{*}_{k})$. Once the iteration reaches the maximum number or the optimization converges, we use the current observation $(\mathbf{p}^{\prime},\mathbf{s}^{\prime})$ as the optimized one $(\mathbf{p}^{*},\mathbf{s}^{*})$. Subsequently, we update the global box $\mathcal{G}^{i}_{t}$ with the finally optimized parameters $(\mathbf{p}^{*},\mathbf{s}^{*})$. The proposed random optimization is performed once the number of candidate boxes $\Psi^{i}_{t}$ is larger than a threshold $\tau_{box}$, or new candidates are added.
4 Experiments
@@ 4.1 Experimental Setup
Baselines: We mainly compare our methods to state-of-the-art offline point-cloud based methods: FCAF [34], TR3D [35] and SpatialLM [39]. Since there are no online open-vocabulary 3D detection alternatives, we take online perception methods: EmbodiedSAM [44] and OnlineAnySeg [38] into consideration, which are initially designed for instance segmentation and are post-processed for gravity-aware 3D bounding box detection following [7].
Metrics: We use the standard metrics for 3D object detection following [30, 1], including Average Precision (AP) at different IoU thresholds. For CA-1M and ScanNetV2, we report class-agnostic $AP$ at IoU thresholds of 0.15, 0.25, and 0.5. ScanNetV2 only provides ground-truth axis-aligned bounding boxes, which requires results of all methods to be axis-aligned for fair comparison. The running efficiency of the system FPS and GPU memory usage are considered as well.
@@
Figure 5: Gallery of 3D object detection on CA-1M and ScanNetV2. Our method is reconstruction-free, and the mesh is semi-transparent. Our method can accurately detect both the common objects and the objects that are not included in most closed-set datasets. Offline methods show poor generalization on fine-grained objects, and OnlineAnySeg (represented as the online method) is slightly better than offline methods but still struggles to detect fine-grained objects. Best viewed on screen.
Figure 6: Visualization of online 3D object detections on 42898849 of CA-1M. The first row shows the detected bounding boxes alongside the camera trajectory (gray), with boxes colored from blue to red to indicate temporal order. The second row displays the RGB images with projected 3D boxes and corresponding depth images at different timestamps.
@@ 4.2 Quantitative Results
The quantitative comparison of our method and state-of-the-art baselines on CA-1M and ScanNetV2 is shown in Table 1. All sequences of the CA-1M validation set and 100 uniformly sampled sequences of the ScanNetV2 validation set are considered. Class-agnostic detection is adopted, since it is easier to obtain the categories of objects with the help of visual foundation models. Our method achieves the best performance on both datasets among the online methods, especially on CA-1M, where our method outperforms the second-best method by a large margin ($22.02$ improvement in $AP_{15}$). Point-based methods [34, 35, 39] demonstrate poor generalization on CA-1M, which is significantly lower than online alternatives in terms of $AP_{15}$ to $AP_{50}$. The main reason is that point-based methods are trained on closed-set datasets, which focus on room-level object detection and typically ignore other ordinary and daily objects.
@@
Online methods like EmbodiedSAM [44] and OnlineAnySeg [38] are designed for instance segmentation, and the results are transformed into bounding boxes for evaluation, following [20]. Details about the transformation are demonstrated in the supplementary materials. Online methods aim to detect objects with streaming RGB-D inputs, which overall show better performance than offline methods. This may be attributed to that all online methods in Table 1 can efficiently capture the rich and useful contexts in images with VFMs for more comprehensive scene understanding.
@@
Point-based method [34, 35] outperforms the others on ScanNetV2. These methods are trained on closed-set datasets and struggle to detect objects of other common categories, which is discussed in Section 4.3. $AP_{15}$ of our method is over 29 on both CA-1M and ScanNetV2, demonstrating the robustness of our method. Moreover, the significant improvement of $AP_{25}$ (20.32 and 6.3) on CA-1M and ScanNetV2 compared to the best online method illustrates that our method can detect objects effectively and accurately without the need for dense reconstruction, which is a significant advantage over other methods. Note that the FPS of OnlineAnySeg [38] is only for the fusion stage, and the segmentation of images is offline, which is time-consuming. Our method is over 20 FPS on average, which can detect and obtain the open-vocabulary semantic features or categories in real time.
4.3 Qualitative Results
Figure 6 presents the visualization of online scanning of BoxFusion. The first row is the detected object boxes, which are colored from blue to red, indicating the temporal order. The second row displays the RGB image with 3D boxes projected into the image planes and the depth image, corresponding to the gray camera trajectory in the first row. Results show that our method can detect fine-grained objects in real time accurately and robustly.
As shown in Figure 5, we compare the detections of our methods to other state-of-the-art baselines. The background of our method is semi-transparent, since we do not rely on reconstruction for detection. In the first two rows, the comparison of two sequences of the CA-1M dataset is displayed, where the ground-truth is more fine-grained than ScanNetV2. Our method stands out as the best with detailed and accurate method. FCAF [34] is trained on closed-set datasets, which can detect the common objects (e.g., sofas and tables) accurately but struggle to detect the fine-grained objects (e.g., newspapers on the coffee table). SpatialLM [39] can only detect the big furniture accurately, and there are many missing objects in the detection. OnlineAnySeg [38] can successfully detect both the common objects and previously unseen objects thanks to the VFM. However, it is hard for OnlineAnySeg to detect the small but important objects (e.g., books and keys at the upper right of the first row). This can be attributed to the merging strategies that rely on point clouds for spatial overlaps, which may fail due to noise or errors in the reconstruction. In contrast, our method is reconstruction-free and leverages the rich context in multi-view images to detect and fuse the bounding boxes of 3D objects, leading to fine-grained and accurate 3D object detection.
The detailed qualitative comparison is demonstrated in Figure 7. Our method preserves the most accurate and complete 3D detection, while other methods indicate inferior performance. The while arrows highlight the comparison. For example, our method can accurately detect the box of stool that is near the wall and partially occluded (the first row). The wall switch (the second row) is successfully detected and multi-view consistent for our method, while other methods fail to detect it.
As shown in Figure 9, our method is open-vocabulary and supports object retrieval with user-specific text prompts. For example, if the prompt indicates it needs to charge the phone (the first row), the power outlet on the wall marked by the red rectangle is highlighted to correspond to the prompt. This indicates that our method can be easily adapted to many downstream embodied tasks.
Figure 7: Detailed qualitative comparison on ScanNetV2 and CA-1M. Our method achieves comprehensive detection of both common objects and fine-grained entities (e.g., the occluded stool, the wall switch), while existing approaches exhibit limited perception accuracy and robustness. White arrows highlight the comparison.
4.4 Ablation Studies
@@
  • SA: Without the SA module, the $AP_{15}$ drops by 18.48, which indicates that the spatial association is important for global 3D object detection. It can provide spatially overlapping details to associate box proposals from different views.
  • MVF: The proposed multi-view box fusion is utilized to generate consistent 3D bounding boxes based on the association modules. There is significant improvement in $AP_{15}$ and $AP_{25}$ (2.09 and 0.51 improvements) compared to not using the MVF module, which validates the effectiveness.
@@
Method $AP_{15}$ $\uparrow$ $AP_{25}$ $\uparrow$ FPS $\uparrow$ GPU mem. $\downarrow$
@@
Figure 8 shows the comparison of different fusion strategies on the ablation sets in Table 2 on CA-1M. The baseline is the strategy that uses the box with the highest score to represent the candidate boxes. The improvements of different strategies compared to the baseline are presented. Additionally, gradient-based fusion strategies using the IoU as the objective loss are overall inferior to the baseline and are not included in the comparison. Our fusion strategy based on particle filtering and random optimization is overall superior to the simple strategy, like averaging. Notably, the improvement of our method compared to averaging is 0.72 in $AP_{50}$, which validates the effectiveness of our method.
Figure 8: Ablation studies of different fusion strategies. Improvements compared to the baseline that uses the box with the highest score as the representative are shown. Random optimization of our method is overall superior to other simple strategies.